How do I solve this equation?
sqrt9x^2 = x + 6

Answer by mananth(16949)

sqrt9x^2 = x + 6
square the equation
9x^2=(x+6)^2
9x^2=x^2+12x+36
8x^2-12x-36=0
Divide the equation by 4
4x^2-3x-9=0
X1,X2 = -b+/- Sqrtb^2-4ac /2a
a=4, b=-3, c= -9
x1= 3+sqrt(9+144) / 8
x1= 1.921
Similarly x2=3-sqrt(9+144) / 8
x2=-1.17